OpenStudy (cutiecomittee123):

How do I find the coordinates of the vertex of a parabola? Also How do I find the Foci?

OpenStudy (coconutjj):

Complete The Square Method.

OpenStudy (hwyl):

the vertex is the highest or the lowest point of a parabola in the form \(f(x) = a(x-h) + k \) (h, k) is the vertex on condition that a is not equal to zero

OpenStudy (loser66):

What is the parabola form? particular problem, please

OpenStudy (cutiecomittee123):

OMG wow I meant to say hyperbola

OpenStudy (coconutjj):

well that's a different story...

OpenStudy (cutiecomittee123):

and the equation given is x^2/16 - y^2/25 =1

OpenStudy (coconutjj):

I don't think there is a vertex...

OpenStudy (cutiecomittee123):

well there are vertices

OpenStudy (cutiecomittee123):

because with a hyperbola there are basically two parabolas

OpenStudy (coconutjj):

dy/dx*

OpenStudy (loser66):

YOur hyperbola has the form \(\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1\) where \(a^2 = 16\) and \(b^2 = 25\)

OpenStudy (loser66):

hence, its graph is